Calculate the probabilities using geometric distribution

Round to 3 decimal places.

a. The first success will occur on the 6th trial with p=0.41
P(6)=

b. The first success will occur on the 9th trial with p=0.2
P(9)=

c. The first success will occur on the 12th trial with p=0.09
P(12)=
Answer #1

If X denotes the number of trials to get first success, then X ~ Geometric(p)

The probability mass function of X is

PlX z) = p(1 _ p)-1, x = 0. 1. 2.3.

a) X ~ Geometric(p = 0.41)

The probability that the first success will occur on the 6th trial

P(X = 6) = 0.41(1-0.41)6-1-0.02931 19

b) X ~ Geometric(p = 0.2)

The probability that the first success will occur on the 9th trial

P(X = 9) = 0.2(1-02)9-1-0.0335544

c)  X ~ Geometric(p = 0.09)

The probability that the first success will occur on the 12th trial

P(X 12) = 0.09(1-0.09)12-1-0.03 18932